The length of a rectangle is 6 meters more than the width. If the perimeter is 128 meters, what are the length and the width?
Answer by Cromlix(4381)

Hi there,
Make Width = x
so, Length = x + 6
Perimeter = 2 x Length + 2 x Width
128 = 2(x + 6) + 2x
Remove brackets
128 = 2x + 12 + 2x
Collect like terms
128 - 12 = 4x
4x = 116
x = 29
Length = 35m
Width = 29m
Hope this helps:-)